OpenStudy (anonymous):

Need help finding the general solution to a differential equation! x(x+y)y'+y(3x+y)=0

OpenStudy (anonymous):

u have to change the variable y to ux so y=ux

OpenStudy (anonymous):

then y'=u'x+u put the new variable instead y to the equation.

OpenStudy (anonymous):

whats u'?

OpenStudy (anonymous):

it is clear that u=y/x

OpenStudy (anonymous):

u prime

OpenStudy (anonymous):

derivative of u

OpenStudy (anonymous):

oh yes for the relationship between them

OpenStudy (anonymous):

this is a way for homogeneous diff equ

OpenStudy (anonymous):

I just don't know how to set it up, to be honest

OpenStudy (anonymous):

ok ill help u

OpenStudy (anonymous):

would it just be divide x(x+y) and sub

OpenStudy (anonymous):

no just put y as a function of u and x

OpenStudy (anonymous):

lets show

OpenStudy (anonymous):

x(x+ux)(u'x+u)=ux(3x+ux)

OpenStudy (anonymous):

then omit the x and solve

OpenStudy (anonymous):

so...(x+ux)(u'x+u)=ux(3x+ux)->>>(1+u)(u'x+u)=u(3+u) is this right??

OpenStudy (anonymous):

pretty good.

OpenStudy (anonymous):

because that's what I did and got the wrong answer

OpenStudy (anonymous):

ok let us check

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

(1+u)u'x=3u+u2-u2-u=2u____>(1+u)u'/u=2/x

OpenStudy (anonymous):

lnu+u=2lnx+k

OpenStudy (anonymous):

or ln(y/x)-ln(x2)=k-y/x

OpenStudy (anonymous):

or y/x3=Cexp(y/x)

OpenStudy (anonymous):

or y= C.x^3.exp(y/x) that Cis a constant

OpenStudy (anonymous):

the answer in the back of the book says: x^2(2xy+y^2)=C

OpenStudy (anonymous):

No the answer isn't right. I checked again.

OpenStudy (anonymous):

I guess you checked the answer in a wrong number.

OpenStudy (anonymous):

check it out.

OpenStudy (anonymous):

No, 100% positive. I got the same thing as you the first time I did it, which is why I was confused. But I can assure you, I am looking at the right answers.

OpenStudy (anonymous):

im sure that the answer i gave u is right.

OpenStudy (anonymous):

I don't know what to say because the answer in the back of the book says something different.

OpenStudy (anonymous):

maybe it can be simplified further?

OpenStudy (usukidoll):

which book are you using?

OpenStudy (usukidoll):

oh it's writing in implicit form. *I got a strange feeling that I"ve this problem before*

OpenStudy (anonymous):

Im using differential equations & linear algebra by Edwards and Penney

OpenStudy (anonymous):

terrible book lol

OpenStudy (usukidoll):

dude try differential equations by bear. VERY HORRIBLE

OpenStudy (anonymous):

well its not really up to me. I have to have it for my math class

OpenStudy (usukidoll):

but there are other ebook s online right?

OpenStudy (anonymous):

yeah, I would normally just watch youtube vids, but I get what to do here. Im just not getting the same answer as the book

OpenStudy (anonymous):

this is the only one that is giving me trouble

OpenStudy (usukidoll):

the book wrote it in implicit form... I don't think your answer is wrong.. just have to ummm rewrite it in a certain way

OpenStudy (anonymous):

the thing is no matter how I write it I will still have an Ln which will lead to an exp. which is not in the answer

OpenStudy (usukidoll):

must be a conversion.. like e^(ln x) would be an x

OpenStudy (anonymous):

the only thing I can think of is to divide both sides by the y in y= C.x^3.exp(y/x) and the ln both sides

OpenStudy (anonymous):

but I think im just going to ask my teacher because I feel like the answer is right

OpenStudy (anonymous):

Thanks but I need to know how to do it. :/

OpenStudy (anonymous):

You can register at wolframapha and get two step by step solution each day

OpenStudy (anonymous):

Here is a solution that gives the answer in the book. I am attached my scanned solution @UsukiDoll @Thefaceless

OpenStudy (anonymous):

@ganeshie8